# Auditory–Tactile Congruence for Synthesis of Adaptive Pain Expressions in RoboPatients

### v1

@@ 5.4 Poststudy Questionaire
We presented 7 Likert scale questions, followed up by an open ended question to obtain any additional thoughts on the experiment.
As the first question we asked if any part of the experiment particularly challenging or difficult with an option to pick four options ranging from ‘very difficult’ to ‘very easy’. We conducted a two-tailed t-test of two samples with unequal variance on the male and female preferences for this question, and the two groups seem to have no significant difference in their preference based on gender (p=0.533 $>$ 0.05). Similarly, the second question was to rate how comfortable they felt while performing the palpations. We did not observe any gender specific preferences in the answers for this either (p=0.146 $>$ 0.05).
We further asked which attribute the participant paid attention to first: facial expression or pain sound, or both when they were performing palpations. We observed that 80% of men paid attention to pain sound and 10% to facial expressions first, and only 10% for both, while only 60% of women paid attention to both and 40% to only pain sound. Overall 70% of the participants paid attention to just the pain sound when palpating the robopatient.
The next question was to investigate if participants were able to distinguish between different levels of pain intensity based on sound and 80% recorded either “distinguishable” or “very distinguishable”. Similar to previous responses, there was no significant difference between the two genders in perceiving the pain intensity in the form of sound (p=0.777 $>$ 0.05). As the next question, we asked if the participant feel any empathy towards the person you were palpating based on the pain sounds and only 25% of the participants responded ”yes” or ”very much”. There was no significant difference in the preferences based on gender for this question either (p=0.423 $>$ 0.05). We then asked if they find it easier or harder to focus on the task as the experiment progressed to gain insights on the design of the experiment. 45% of the participants responded that the experiment was getting easier over time while 55% had an opposite impression. Based on the two-tailed t-test performed on the responses to this question, there was no gender-based significance in these preferences (p = 0.343 $>$ 0.05). Finally, we asked if the participant recollected any memories of their own experience about an encounter with a doctor in the past during the experiment, and 30% responded “yes” to this question and choices were equally distributed among the two genders.
@@
Under additional comments from the participants, were suggestions to incorporate the speed of palpation to map appropriate pain sounds and include a more diverse set of pain sounds. Furthermore, some commented that it was easier to distinguish female pain sounds compared to that of a male, and it would have been useful to know more about a certain injury and the type of pain to have more resemblance to the scenario.
@@ 5.5 Limitations, Implications and Future Work
Furthermore, the pain-to-sound mapping was limited to three levels, making it difficult for participants to differentiate feedback at maximum pain intensity. Additionally, the small sample size of naive participants may have reduced the statistical power, resulting in non-significant outcomes. This study is presented as a pilot to demonstrate the integration of vocal pain feedback into robopatients for abdominal palpation training. We plan to expand our investigations with a larger number of participants from diverse genders and ethnic backgrounds, and participants with different palpation experience levels to address these limitations. In addition we intend to expand the experiments to medical professionals. In addition, time scales of pain sound is also important to consider, for instance, higher pitch and louder pain cries are associated with a sense of urgency [26, 44]. Hence, these properties will be incorporated into future pain expression generation.
@@ 6 Conclusions
This study examined how acoustic features—specifically amplitude and pitch—interact with user-applied force thresholds to shape perceptual judgments of congruence in a palpation training context. Through a comprehensive analysis we observed that participant agreement with auditory feedback was strongly modulated by auditory characteristics rather than the identity of the pain sound or the magnitude of tactile force alone. Key findings from ANOVA and post-hoc tests revealed that amplitude and pitch are central to shaping perceived pain congruence in robot-mediated palpation, with louder and mid-range pitch values more often eliciting agreement. Force threshold also had a significant effect, particularly with ‘strongly agree’ responses being linked to higher forces. However, pain sound identity (male or female voice) showed no statistically significant influence, indicating that speaker characteristics were not perceptually salient in this context. Cluster analysis in PCA space identified two distinct perceptual groupings corresponding to different combinations of pitch and amplitude. These clusters aligned significantly with participants’ choices, suggesting that perceived congruence emerges from nuanced auditory-tactile mappings rather than isolated parameter changes. Importantly, Random Forest classification confirmed pitch as the most dominant predictor of participant clustering, followed by amplitude. Force threshold and pain sound had negligible effects on prediction performance. This suggests that, although force plays a role in increasing confidence at higher thresholds, it is not the primary determinant of perceived congruence when paired with auditory feedback. As a whole, our results point to a perceptual framework where pitch and amplitude serve as core dimensions in shaping the realism and acceptability of robotic auditory feedback in simulated palpation tasks. This has direct implications for the design of robopatient systems, emphasizing the need for fine-grained acoustic modulation over simplistic force-to-sound mappings or reliance on speaker identity. For training scenarios that rely on simulating affective states such as pain, these insights can guide the development of more intuitive, convincing, and effective human-robot interaction systems.
@@ Acknowledgements
The authors are thankful to the participants who generously volunteered for this study. Each participant received a £10 Amazon voucher as a token of appreciation. This work was partially supported by the School of Engineering and Materials Science (SEMS) at Queen Mary University of London, Start-up funding.
This work is supported by the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101034337.

### v2

@@ 5.4 Poststudy Questionaire
We presented 7 Likert scale questions, followed up by an open-ended question to obtain any additional thoughts on the experiment.
As the first question, we asked if any part of the experiment was particularly challenging or difficult, with an option to pick four options ranging from ‘very difficult’ to ‘very easy’. We conducted a two-tailed t-test of two samples with unequal variance on the male and female preferences for this question, and the two groups seem to have no significant difference in their preference based on gender (p=0.533 >> 0.05). Similarly, the second question was to rate how comfortable they felt while performing the palpations. We did not observe any gender specific preferences in the answers for this either (p=0.146 >> 0.05).
We further asked which attribute the participant paid attention to first: facial expression or pain sound, or both, when they were performing palpations. We observed that 80% of men paid attention to pain sound and 10% to facial expressions first, and only 10% for both, while only 60% of women paid attention to both and 40% to only pain sound. Overall 70% of the participants paid attention to just the pain sound when palpating the robopatient.
The next question was to investigate if participants were able to distinguish between different levels of pain intensity based on sound, and 80% recorded either “distinguishable” or “very distinguishable”. Similar to previous responses, there was no significant difference between the two genders in perceiving the pain intensity in the form of sound (p=0.777 >> 0.05). As the next question, we asked if the participant feel any empathy towards the person you were palpating based on the pain sounds, and only 25% of the participants responded ”yes” or ”very much”. There was no significant difference in the preferences based on gender for this question either (p=0.423 >> 0.05). We then asked if they found it easier or harder to focus on the task as the experiment progressed to gain insights into the design of the experiment. 45% of the participants responded that the experiment was getting easier over time while 55% had an opposite impression. Based on the two-tailed t-test performed on the responses to this question, there was no gender-based significance in these preferences (p = 0.343 >> 0.05). Finally, we asked if the participant recollected any memories of their own experience about an encounter with a doctor in the past during the experiment, and 30% responded “yes” to this question, and choices were equally distributed among the two genders.
@@
Under additional comments from the participants, were suggestions to incorporate the speed of palpation to map appropriate pain sounds and include a more diverse set of pain sounds. Furthermore, some commented that it was easier to distinguish female pain sounds compared to those of a male, and it would have been useful to know more about a certain injury and the type of pain to have more resemblance to the scenario.
@@ 5.5 Limitations, Implications and Future Work
Furthermore, the pain-to-sound mapping was limited to three levels, making it difficult for participants to differentiate feedback at maximum pain intensity. Additionally, the small sample size of naive participants may have reduced the statistical power, resulting in non-significant outcomes. This study is presented as a pilot to demonstrate the integration of vocal pain feedback into robopatients for abdominal palpation training. We plan to expand our investigations with a larger number of participants from diverse genders and ethnic backgrounds, and participants with different palpation experience levels, to address these limitations. In addition, we intend to expand the experiments to medical professionals. Furthermore, time scales of pain sound is also important to consider, for instance, higher pitch and louder pain cries are associated with a sense of urgency [36, 11]. Hence, these properties will be incorporated into future pain expression generation.
@@
Alongside pain-to-sound mapping, we will also improve the physical appearance of the robopatient to resemble a real patient lying on a hospital bed. Mainly, we will improve the design of the abdominal phantom by adding details such as the navel and abdominal muscles. In addition, better integration with MorphFace will improve the participant’s perception and empathy during palpation tasks and foster a more realistic interaction with the robopatient.
@@ 6 Conclusions
This study examined how acoustic features, specifically amplitude and pitch, interact with user-applied force thresholds to shape perceptual judgments of congruence in a palpation training context. Through a comprehensive analysis, we observed that participant agreement was primarily modulated by the intensity of the interaction, specifically the combination of auditory amplitude and tactile force, rather than the identity of the pain sound. Key findings from ANOVA and post-hoc tests revealed that amplitude and force threshold are the central drivers of perceived pain congruence in robot-mediated palpation, with louder sounds and higher force levels significantly linked to ‘strongly agree’ responses. Pitch played a significant but secondary role, while pain sound identity (male or female voice) showed no statistically significant influence, indicating that speaker characteristics were not perceptually salient in this context.
Visualization in PCA space supported these statistical findings, revealing that participant choices form a gradient based on stimulus magnitude rather than distinct, isolated clusters. The data indicates that users rely on a magnitude-based heuristic to judge realism: high-intensity cues (high force paired with high amplitude) were most consistently linked to strong agreement. Conversely, mismatches in intensity (e.g., high force with low sound) led to disagreement.
The auditory pain expressions of robopatients can be optimized by altering the pitch and amplitude dynamically based on the palpation forces, with the specific audio having minimal significance. In our future work, we plan to conduct a large-scale study to further investigate and reinforce the findings of this work, and improve the robopatient to test in a clinical training environment.
As a whole, our results point to a perceptual framework where pitch and amplitude serve as core dimensions in shaping the realism and acceptability of robotic auditory feedback in simulated palpation tasks. This has direct implications for the design of robopatient systems, emphasizing the need for fine-grained acoustic modulation over simplistic force-to-sound mappings or reliance on speaker identity. For training scenarios that rely on simulating affective states such as pain, these insights can guide the development of more intuitive, convincing, and effective human-robot interaction systems.
@@ Acknowledgements
The authors are thankful to the participants who generously volunteered for this study. Each participant received a £10 Amazon voucher as a token of appreciation. This work was partially supported by the School of Engineering and Materials Science (SEMS) at Queen Mary University of London start-up funding and the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101034337.